OpenStudy (anonymous):

LAST QUESTION ! I PROMISE ! PLEASE HELP ! What is the sum of the geometric sequence 4, 12, 36 . if there are 9 terms?

OpenStudy (darkprince14):

find r then use the formula in finding the sum of geometric series... r stands for the ratio of the first and the next term...

OpenStudy (anonymous):

a1= 4 r=12/4=3

OpenStudy (anonymous):

\[\Sigma = a\frac{1-r^n}{1-r}\]

OpenStudy (darkprince14):

\[a \left( \frac{ 1-r ^{n} }{ 1-r } \right)\] a is the first term r is the "common ratio" between terms n is the number of terms

OpenStudy (darkprince14):

geometric sequence.... sorry for using the term 'series'

OpenStudy (anonymous):

No, series is right

OpenStudy (anonymous):

hahahnp ! now what am I supposed to do with that equation

OpenStudy (darkprince14):

substitute the values

OpenStudy (anonymous):

You already know \(a\) and \(r\)/ \(n\) is the number of terms, 9.

OpenStudy (anonymous):